HETAV PATEL

Grade 11,

A stone is drooped from the top of a tower travels 25 m in last second of its motion. If [g=10], height of the tower is

A stone is drooped from the top of a tower travels 25 m in last second of its motion. If [g=10], height of the tower is

Grade:7

3 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,We know distance travelled in nth second = u +(a/2)(2n-1)Here u=0, a=g and n = total time takenGiven distance travelled in last second = 25mSo 25= (10/2)(2n-1)=> 5=2n-1=> n=3 seconds = total time taken.height of tower= (1/2)*g*n^2= (1/2)*10*3*3=45 m.So height of tower is 45 m. [Ans]Thank you,Soumya Ranjan Mohanty.
rahul gurjar
6 years ago
answere ,Dear student,We know distance travelled in nth second = u +(a/2)(2n-1)Here u=0, a=g and n = total time takenGiven distance travelled in last second = 25mSo 25= (10/2)(2n-1)=> 5=2n-1=> n=3 seconds = total time taken.height of tower= (1/2)*g*n^2= (1/2)*10*3*3=45 m.So height of tower is 45 m. [Ans]Thank you,Soumya Ranjan Mohanty.
sahil
5 years ago
the
the answer is
 ,Dear student,We know distance travelled in nth second = u +(a/2)(2n-1)Here u=0, a=g and n = total time takenGiven distance travelled in last second = 25mSo 25= (10/2)(2n-1)=> 5=2n-1=> n=3 seconds = total time taken.height of tower= (1/2)*g*n^2= (1/2)*10*3*3=45 m.So height of tower is 45 m. [Ans]Thank you,Soumya Ranjan Mohanty.
is